OpenStudy (anonymous):

Write an equation of a line with a slope of 0 and passing through the point (5,4).

OpenStudy (anonymous):

From your last question we know that any line with a slope of zero is represented by the equation y = c where c is a constant Now we were also given a point of the graph that tells us that when x is 5 y is 4 and we can plug those two values into our formula for the general equation of a line with a slope of zero which gives y = c 4 = c We now take the value of c and plug that back into the general formula y = 4 is the equation of the line Another way to get the general formula is to sub in the value of zero for the slope (m) to the general line equation y = m*x + c y = 0*x + c y = c

OpenStudy (anonymous):

ok that made sense. Can I ask you three more that are similar to this one?

OpenStudy (anonymous):

i think i already have the enxt one, so it would only be 2 more questions

OpenStudy (anonymous):

yea same thing though also make sure you are closing each question as we finish them

OpenStudy (anonymous):

ok.:)